An aqueous solution of perchloric acid is standardized by titration with a 0.108 M solution of barium hydroxide.

If 24.8 mL of base are required to neutralize 14.4 mL of the acid, what is the molarity of the perchloric acid solution?

__________M perchloric acid
Answer #1

Balanced chemical equation is:

Ba(OH)2 + 2 HClO4 ---> Ba(ClO4)2 + 2 H2O

Here:

M(Ba(OH)2)=0.108 M

V(Ba(OH)2)=24.8 mL

V(HClO4)=14.4 mL

According to balanced reaction:

2*number of mol of Ba(OH)2 =1*number of mol of HClO4

2*M(Ba(OH)2)*V(Ba(OH)2) =1*M(HClO4)*V(HClO4)

2*0.108*24.8 = 1*M(HClO4)*14.4

M(HClO4) = 0.372 M

Answer: 0.372 M